OpenStudy (anonymous):

Can i have some help with this? √75 + √3

OpenStudy (anonymous):

does it help to know that \(75=25\times 3\)?

OpenStudy (anonymous):

use the website mathway!

OpenStudy (anonymous):

so \[\sqrt{75}=\sqrt{25\times 3}=\sqrt{25}\times \sqrt3\]

OpenStudy (anonymous):

you can enter in the equation and it gives you the answer

OpenStudy (anonymous):

you end up with \[5\sqrt3+\sqrt3\] combine like terms to finish

OpenStudy (anonymous):

Thanks @ihenandezzi but i'm trying to learn how to do the problem, besides i have the link :)

OpenStudy (anonymous):

@satellite73 wouldn't the like term make 6 or do i just write it out and don't add them?

OpenStudy (anonymous):

Anyone?

OpenStudy (solomonzelman):

ok, we got to the point of \(5\sqrt{3}+\sqrt{3}\).

OpenStudy (solomonzelman):

Do you know what is 5x+x ?

OpenStudy (anonymous):

7?

OpenStudy (solomonzelman):

5x+x=5x+1x=6x ---------------

OpenStudy (solomonzelman):

is that correct?

OpenStudy (anonymous):

Yeah

OpenStudy (blake4560):

yes, 5x+x=6x

OpenStudy (solomonzelman):

5x+x=6x and same way, it follows that 5√3 + √3 = ?

OpenStudy (anonymous):

5√6

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

you are not adding inside the square root

OpenStudy (anonymous):

So i don't add the 3's

OpenStudy (solomonzelman):

you are adding the "square-root-of-3" five times, PLUS "square-root-of-3" one time.

OpenStudy (blake4560):

isn't it \[5\sqrt{3}\]

OpenStudy (solomonzelman):

\(5\sqrt{3}+\sqrt{3}=5\sqrt{3}+1\sqrt{3}=?\)

OpenStudy (anonymous):

I don't know. guessing 2?

OpenStudy (solomonzelman):

5+1=?

OpenStudy (anonymous):

6

OpenStudy (solomonzelman):

5a+1a=?

OpenStudy (blake4560):

6a

OpenStudy (solomonzelman):

5(something) + 1(something) = ?

OpenStudy (anonymous):

Depends on the something.

OpenStudy (solomonzelman):

6(something)

OpenStudy (solomonzelman):

5(something) + 1(something) = 6(something) (as long as we are talking about the same something)

OpenStudy (solomonzelman):

5√3 + 1√3 = ?

OpenStudy (anonymous):

Think i said 6..

OpenStudy (solomonzelman):

5√3 + 1√3 = 6 ?

OpenStudy (anonymous):

5√6

OpenStudy (anonymous):

1√3

OpenStudy (solomonzelman):

5 + 1 = 6 5x + 1x = 6x 5 bees + 1 bee = 6 bees 5a + 1a = 6a 5√3 + 1√3 = 6√3 compare these

OpenStudy (anonymous):

6√6

OpenStudy (solomonzelman):

why?

OpenStudy (anonymous):

3+3 5+1 6,6

OpenStudy (solomonzelman):

you are adding the number of times of which you have a √3. you have five square roots of 3, and then you add one more square root of 3. how many square roots of 3 do you have together?

OpenStudy (anonymous):

6<---

OpenStudy (solomonzelman):

yes, so 5√3 + 1√3 = ?

OpenStudy (anonymous):

Do we add the 3's?

OpenStudy (anonymous):

if so that would be 6√6

OpenStudy (solomonzelman):

we don't add the square root inside, just 5√3+1√3=(5+1)√3 = 6√

OpenStudy (solomonzelman):

6√3

OpenStudy (anonymous):

Okay.